If the revenue R and demand Q are related by R = Q(50 – 2Q) find the marginal revenue function. What is the marginal revenue when Q = 10?

(Answer. R’ = 50 – 4Q, so when Q = 10, marginal revenue is 10.)
Answer #1

Since; marginal Revenue ,AR= d (total Relowe) dQ TMR = a(R) da to differential equation of MR here R= 500-202 MR= d(500 - 2